OpenStudy (141823):

The probability of a randomly chosen boy playing basketball is 0.30. The chance that a boy plays both basketball and football is 0.05. The chance that a boy plays football is 0.25. What is the probability that a randomly chosen boy plays basketball or football?

OpenStudy (jango_in_dtown):

Let B denote the event, a boy plays basketball and F denote the event a boy plays foot ball. Then by question P(B)=0.30, P(F)=0.25 and P(B.F)=0.05 Required probability =P(B+F)=P(B)+P(F)-P(B.F)=0.30+0.25-0.05

OpenStudy (jango_in_dtown):

@141823

OpenStudy (alivejeremy):

ก็็็็็็็็็็็็็ʕ•͡ᴥ•ʔ ก้้้้้้้้้้้

OpenStudy (jadeishere):

@jango_IN_DTOWN I think you should explain each individual step as a whole. People who are just learning this stuff will find it very confusing unless thoroughly explained :) But wonderful job helping!

OpenStudy (141823):

Sorry I was afk! Thank you for your help, so it looks like for this problem we would use the overlapping formula, P(B+F)=P(B)+P(F)-P(BnF) Thank you for that help, I was just not sure which formula I would need to follow when they use the term "or". In the end our answer would be .5?

OpenStudy (jango_in_dtown):

right